OpenStudy (butterflyhope):

Urgent *WILL MEDAL!*: Which function represents exponential growth? y=9/7(2/3)^x y=4/5(4/5)^x y=2.3(3/2)^x y=4.1(5/6)^x Thanks! :)

jimthompson5910 (jim_thompson5910):

Rule: \[\Large y = a*b^x\] has exponential growth only if b > 1

jimthompson5910 (jim_thompson5910):

For example, \[\Large y = 2*1.7^x\] represents growth because b = 1.7 which is larger than 1

OpenStudy (butterflyhope):

Ohhh I get it but none of them are greater than 1??

jimthompson5910 (jim_thompson5910):

convert each of the fractions to decimal form that should help

jimthompson5910 (jim_thompson5910):

ex: 9/7 = 1.29 2/3 = 0.67 y=9/7(2/3)^x turns into y = 1.29*0.67^x the base for x is 0.67 which is NOT greater than 1, so it is NOT a growth function. It's actually decaying So A is out

OpenStudy (butterflyhope):

It's C because 3/2 = 1.5?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (butterflyhope):

Thank you :)

jimthompson5910 (jim_thompson5910):

1.5 = 1+r solve for r to get r = 0.5 since r = 0.5 = 50%, we have 50% growth here

jimthompson5910 (jim_thompson5910):

you're welcome